OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y =1/2 x - 8. y =1/2 x + 7 y =1/2 x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (ajprincess):

Slopes are equal for parallel lines. So slope of the required line=1/2 \(y-y_1=m(x-x_1)\) Here m=1/2(slope) \(x_1\)=-2 \(y_1\)=6 Plug in these values in the above eqn and solve for y. Can u do it @carachronic?

OpenStudy (anonymous):

its the first one y =1/2 x + 7

OpenStudy (anonymous):

Best response?

OpenStudy (anonymous):

y=1/2x-8 6=1/2(-2)+7 6=-1+7 6=6

OpenStudy (anonymous):

ohh kinda hard i see though kinda (:

OpenStudy (anonymous):

y = 1/2 x + 7 :) since its parallel, the slope stays the same: 1/2 then.... since it passes through (-2, 6), therefore: y-6 = 1/2(x+2) y - 6 = 1/2x +1 y = 1/2 x +7